Abardeen Corporation borrowed $85,000 from the bank on October 1, Year 1. The note had an 6 percent annual rate of interest and matured on March 31, Year 2. Interest and principal were paid in cash on the maturity date.

Required
a. What amount of cash did Abardeen pay for interest in Year 1?
b. What amount of interest expense was recognized on the Year 1 income statement? (Do not round intermediate calculations. Round your answer to the nearest dollar amount.)
c. What amount of total liabilities was reported on the December 31, Year 1, balance sheet? (Do not round intermediate calculations. Round your answer to the nearest dollar amount.)
d. What total amount of cash was paid to the bank on March 31, Year 2, for principal and interest?
e. What amount of interest expense was reported on the Year 2 income statement? (Do not round intermediate calculations. Round your answer to the nearest dollar amount.)
Answer #1
a
Cash paid for interest in Year 1 0
b
Interest expense Year 1 1275 =85000*6%*3/12
c
Total liabilities December 31, Year 1 86275 =85000+1275
d
Cash paid for principal and interest 87550 =85000+1275+1275
e
Interest expense Year 2 1275 =85000*6%*3/12